# LayerSync: Self-aligning Intermediate Layers

### v1

@@ Appendix C LayerSync - Algorithm
Appendix D FLOP comparison
We compare the computational complexity of Dispersive Loss, which computes pairwise distances, with LayerSync in Table 9. LayerSync is more efficient in terms of computational complexity as the pairwise comparisons in Dispersive Loss result in a quadratic cost with respect to batch size.
@@ Appendix D FLOP comparison
Appendix E Image generation experimental details
We use a node of 4 GH200 GPUs and a batch size of 256. The details of hyper parameters and sampler are provided in Tables 10 and 11.
@@ Appendix E Image generation experimental details
Appendix F Human motion generation experimental details
@@
We rely on HumanML3D dataset  (Guo et al., 2022a) that contains 44,970 motion annotations across 14,646 motion sequences from the AMASS  (Mahmood et al., 2019) and HumanAct12  (Guo et al., 2020) datasets, along with corresponding text descriptions, and is widely used for the task of text-conditional human motion generation. Motions in the HumanML3D dataset follow the skeleton structure of SMPL  (Loper et al., 2015) with 22 joints. Each pose $\mathbf{p}$ in the motion sequence is represented by a vector of size 237,
@@
We use the exact setup as MDM  (Tevet et al., 2023), we train up to 600K iterations using a H100 GPU. We sync block 3 with block 6.
Appendix G Stochastic Interpolants
@@
Appendix H Evaluation metrics details.
H.1 Image
@@
  • Inception Score. Salimans et al. (2016) computes the Kullback–Leibler (KL) divergence  (Kullback & Leibler, 1951) between conditional and marginal label distributions predicted by an Inception network.
  • Percision and Recall. Kynkäänniemi et al. (2019) measures the fraction of generated samples that lie within the support of the real data distribution in feature space. High recall reflects the diversity of generated samples, indicating that the model captures the variability of the real data distribution.
H.2 Audio
@@
H.3 Video
  • FVD: Unterthiner et al. (2018) extends the idea of FID to videos by measuring the distance between real and generated video distributions in a pretrained spatiotemporal feature space. Specifically, it uses embeddings from Carreira & Zisserman (2017), pretrained on large-scale video datasets, to capture both spatial and temporal dynamics.
H.4 Motion
  • FID: Computed in the same way as for images, but using T2M  (Guo et al., 2022b) motion features instead of Inception features.
  • R-Precision: Measure the relevancy of the generated motions to the input prompts.
Appendix I Extended related work
@@
  • ADM (Dhariwal & Nichol, 2021): Builds upon U-Net-based diffusion models by introducing classifier-guided sampling, allowing fine-grained control over the trade-off between generation quality and diversity.
  • VDM++ (Kingma & Gao, 2023): Proposes an adaptive noise schedule that adjusts dynamically during training, improving convergence and sample quality.
@@
  • DiT (Peebles & Xie, 2023): Proposes a pure transformer architecture for diffusion, using AdaLN-zero modules to stabilize training and scale to large model sizes efficiently.
@@
Appendix J Details of SiT model
The architecture of the SiT block is provided in Figure 6 and more details on the model parameters are summarized in Table 12.
@@
Appendix K Attention Maps PCA over layers
@@
Appendix L Qualitative Examples
We provide qualitative examples in Figure 18. The model is trained for 800 on ImageNet dataset  (Deng et al., 2009) and the samples are generated using classifier-free guidance with a scale of 4 and the ODE Heun sampler. Additional qualitative comparisons between the baseline SiT-XL/2, SiT-XL/2 regularized with Dispersive, and SiT-XL/2 regularized with LayerSync trained on ImageNet dataset  (Deng et al., 2009) are shown in Figure 19. All models are trained for 400K iterations and share the same noise, sampler, and number of sampling steps. The samples are generated using ODE Heun sampler and no classifier-free guidance is used. LayerSync improves generation quality without relying on external representation.
Figure 18: Selected samples from the SiT XL/2 with LayerSync on ImageNet $\times256$. We use classifier free guidance with a cfg of 4.0.

### v2

@@ Appendix I Evolution of Block Structure
Appendix K FLOP comparison
We compare the computational complexity of Dispersive Loss, which computes pairwise distances, with LayerSync in Table 17. LayerSync is more efficient in terms of computational complexity, as the pairwise comparisons in Dispersive Loss result in a quadratic cost with respect to batch size.
@@
Appendix L Image generation experimental details
We use a node of 4 GH200 GPUs and a batch size of 256. The details of hyperparameters and sampler are provided in Tables 18 and 19.
@@
Appendix M Human motion generation experimental details
@@
We rely on HumanML3D dataset  (Guo et al., 2022) that contains 44,970 motion annotations across 14,646 motion sequences from the AMASS  (Mahmood et al., 2019) and HumanAct12  (Guo et al., 2020) datasets, along with corresponding text descriptions, and is widely used for the task of text-conditional human motion generation. Motions in the HumanML3D dataset follow the skeleton structure of SMPL  (Loper et al., 2015) with 22 joints. Each pose $\mathbf{p}$ in the motion sequence is represented by a vector of size 237,
@@
We use the exact setup as MDM  (Tevet et al., 2022), we train up to 600K iterations using a H100 GPU. We sync block 3 with block 6.
Appendix N Stochastic Interpolants
@@
Appendix O Evaluation metrics details.
O.1 Image
@@
  • Inception Score. Salimans et al. (2016) computes the Kullback–Leibler (KL) divergence  (Kullback and Leibler, 1951) between conditional and marginal label distributions predicted by an Inception network.
  • Precision and Recall. Kynkäänniemi et al. (2019) measures the fraction of generated samples that lie within the support of the real data distribution in feature space. High recall reflects the diversity of generated samples, indicating that the model captures the variability of the real data distribution.
O.2 Audio
@@
O.3 Video
  • FVD: Unterthiner et al. (2018) extends the idea of FID to videos by measuring the distance between real and generated video distributions in a pretrained spatiotemporal feature space. Specifically, it uses embeddings from Carreira and Zisserman (2017), pretrained on large-scale video datasets, to capture both spatial and temporal dynamics.
O.4 Motion
  • FID: Computed in the same way as for images, but using T2M  (Guo et al., 2022) motion features instead of Inception features.
  • R-Precision: Measures the relevancy of the generated motions to the input prompts.
Appendix P Extended related work
@@
  • ADM (Dhariwal and Nichol, 2021): Builds upon U-Net-based diffusion models by introducing classifier-guided sampling, allowing fine-grained control over the trade-off between generation quality and diversity.
  • VDM++ (Kingma and Gao, 2023): Proposes an adaptive noise schedule that adjusts dynamically during training, improving convergence and sample quality.
@@
  • DiT (Peebles and Xie, 2023): Proposes a pure transformer architecture for diffusion, using AdaLN-zero modules to stabilize training and scale to large model sizes efficiently.
@@
  • D-JEPA (Chen et al., 2024): Integrates Joint-Embedding Predictive Architectures (JEPA) into generative modeling by reframing masked image modeling as a generalized next-token prediction task, utilizing diffusion or flow matching loss to model per-token probability distributions in a continuous space.
  • VAR (Tian et al., 2024): Redefines autoregressive image generation as a coarse-to-fine ”next-scale prediction” process, diverging from standard raster-scan next-token prediction, allowing for faster inference and scaling laws similar to Large Language Models.
  • SRA (Jiang et al., 2025): An EMA-based method that enables diffusion transformers to enhance their own representation learning and generation quality by aligning latent outputs from earlier, noisier layers with those from later, cleaner layers, eliminating the need for external guidance models.
Appendix Q Details of SiT model
The architecture of the SiT block is provided in Figure 10, and more details on the model parameters are summarized in Table 20.
@@
Appendix R Attention Maps PCA over layers
@@
Appendix S Qualitative Examples
We provide qualitative examples in Figure 22. The model is trained for 800 epochs on ImageNet dataset  (Deng et al., 2009) and the samples are generated using classifier-free guidance with a scale of 4 and the ODE Heun sampler.
Additionally, qualitative comparisons between the baseline SiT-XL/2, SiT-XL/2 regularized with Dispersive, and SiT-XL/2 regularized with LayerSync trained on ImageNet dataset are shown in Figure 23. All models are trained for 400K iterations and share the same noise, sampler, and number of sampling steps. The samples are generated using ODE Heun sampler and no classifier-free guidance is used. LayerSync improves generation quality without relying on external representation.
Figure 22: Selected samples from the SiT XL/2 with LayerSync on ImageNet $\times256$. We use classifier-free guidance with a CFG of 4.0.